OpenStudy (anonymous):

what is the slope of the line -x+4y=8

OpenStudy (loser66):

isolate y the number in the front of x is slope. (coefficient of x)

OpenStudy (loser66):

Practice, please. I can check

OpenStudy (anonymous):

ok give a minute

OpenStudy (anonymous):

I got 1/4+2 is that correct

OpenStudy (loser66):

first of all, you must show the function, I mean y =.... to make sure that you know how to get it second, your chain of numbers has no meaning.

OpenStudy (anonymous):

y=1/4x+2 i am trying to find the slope of the line

OpenStudy (loser66):

yeah, you got it, so? the slope is?

OpenStudy (loser66):

the number in the front of x

OpenStudy (anonymous):

The slope is the m in the equation y = mx + b Once you break the equation down you get y = 1/4x + 2. what is the slope?

OpenStudy (anonymous):

1/4 ok I think I understand thanks for all your help!

OpenStudy (anonymous):

Yes 1/4

OpenStudy (loser66):

Bingo, sorry, I don't have medal to give you. I gave it to MDoodler, hihihi

OpenStudy (anonymous):

lol thanks anyway

OpenStudy (loser66):

give him yours.

OpenStudy (loser66):

don't forget, you must write down the whole equation .